A population is normally distributed with a mean of 100 and a standard deviation of 10, for samples of size 25, what is the probability of randomly sampling and getting a sample mean of 103 or more?
Answer #1

let us consider r ~N(100, 102)

and we have given n = 25

we know that x\sim N(\mu,\sigma^{2})

then \bar{x}\sim N(\mu,\frac{\sigma}{\sqrt{n}})

P (> 103) =(103 - 100) P(z > P10/V25

P (> 103) = P(z 1.5)

P (> 103) = 0.0668